# Physical meaning of non-extensive term in Massieu functions

### v1

@@ Acknowledgements.
Eq. (37) can be used to demonstrate that the difference $\Delta\tilde{\Psi}(\vec{Y})$ for the last Legendre transform is a constant.
We obtain the probability of fluctuations, $\Delta\hat{\vec{X}}=\hat{\vec{X}}-\vec{X}$, directly from the probability of $\hat{\vec{X}}$, Eq. (13). For simplicity, we consider here the ensemble where all extensive variables (except volume) have fluctuations; it is described by the first Massieu function, that is, the entropy $\tilde{\Psi}(\hat{\vec{X}})$. Expanding $\tilde{\Psi}(\hat{\vec{X}})$ in a Taylor series up to second order in $\Delta\hat{\vec{X}}$, we have
$\tilde{\Psi}$ $(\vec{X}+\Delta\hat{\vec{X}})$
$=\tilde{\Psi}(\vec{X})+\frac{\partial\tilde{\Psi}(\vec{X})}{\partial\vec{X}}\cdot\Delta\hat{\vec{X}}+\frac{1}{2}\Delta\hat{\vec{X}}\cdot\frac{\partial^{2}\tilde{\Psi}(\vec{X})}{\partial\vec{X}^{2}}\cdot\Delta\hat{\vec{X}}$
$=\tilde{\Psi}(\vec{X})+\vec{Y}\cdot\Delta\hat{\vec{X}}-\frac{1}{2}\Delta\hat{\vec{X}}\cdot C^{-1}\cdot\Delta\hat{\vec{X}},$ (71)
where we used that $\frac{\partial\tilde{\Psi}(\vec{X})}{\partial\vec{X}}=\vec{Y}$ and that the inverse of the fluctuation matrix (or covariance matrix) is $C^{-1}=-\frac{\partial\vec{Y}}{\partial\vec{X}}$. Replacing in (13), we obtain
$P(\Delta\hat{\vec{X}})=e^{-\tilde{\Psi}(\vec{Y})+\tilde{\Psi}(\vec{X})-\vec{Y}\cdot\vec{X}-\frac{1}{2}\Delta\hat{\vec{X}}\cdot C^{-1}\cdot\Delta\hat{\vec{X}}}.$ (72)
Using the Legendre transform $\Psi(\vec{Y})=\Psi(\vec{X})-\vec{Y}\cdot\vec{X}$, and $\Delta\tilde{\Psi}(\vec{X})=\tilde{\Psi}(\vec{X})-\Psi(\vec{X})$ and $\Delta\tilde{\Psi}(\vec{Y})=\tilde{\Psi}(\vec{Y})-\Psi(\vec{Y})$, we have
$P(\Delta\hat{\vec{X}})=e^{-\Delta\tilde{\Psi}(\vec{Y})+\Delta\tilde{\Psi}(\vec{X})}\,e^{-\frac{1}{2}\Delta\hat{\vec{X}}\cdot C^{-1}\cdot\Delta\hat{\vec{X}}},$ (73)
therefore, the factor $e^{-\Delta\tilde{\Psi}(\vec{Y})+\Delta\tilde{\Psi}(\vec{X})}$, a function of the equilibrium parameters $\vec{X}$ and $\vec{Y}$, is the normalization of the multivariate normal distribution given by $(2\pi)^{-r/2}(\det C)^{-1/2}$, where $r$ is the number of variables; see, for example, Mishin (2015) or Sec. 7.C in Reichl (1998); the multivariate Gaussian distribution is a consequence of the central limit theorem, see Ch. V in Khinchin (1949).
@@
$e^{-\Delta\tilde{\Psi}(\vec{Y})+\Delta\tilde{\Psi}(\vec{X})}=(2\pi)^{-r/2}(\det C)^{-1/2},$ (74)
@@
$\Delta\tilde{\Psi}(\vec{X})-\Delta\tilde{\Psi}(\vec{Y})=-\frac{1}{2}\ln(\det C)-\frac{r}{2}\ln(2\pi).$ (75)
Now, using Eq. (37), we have that $\Delta\tilde{\Psi}(\vec{Y})$ is a constant:
$\Delta\tilde{\Psi}(\vec{Y})=c+\frac{r}{2}\ln(2\pi).$ (76)
@@
with $A=Me^{-2\epsilon/T}/\sqrt{N^{2}-M^{2}}$. Using this equation, the TH Massieu function $\Psi(M,\tfrac{1}{T})$ is
$\Psi(M,$ $\tfrac{1}{T})=\underbrace{N_{+}\ln\frac{N}{N_{+}}+N_{-}\ln\frac{N}{N_{-}}}_{\text{ideal part}}+\frac{\epsilon}{T}(2M-N)$
@@
And the mean squared fluctuations $\langle(\Delta\hat{M})^{2}\rangle=-(\frac{\partial^{2}\Psi(M,1/T)}{\partial M^{2}})^{-1}$ are given by

### v2

@@ Acknowledgements.
Eq. (37) can be used to demonstrate that the difference $\Delta\tilde{S}[\vec{Y}]$ for the last Legendre transform is a constant.
We obtain the probability of fluctuations, $\Delta\hat{\vec{X}}=\hat{\vec{X}}-\vec{X}$, directly from the probability of $\hat{\vec{X}}$, Eq. (13). For simplicity, we consider here the ensemble where all extensive variables (except volume) have fluctuations; it is described by the first Massieu function, that is, the entropy $\tilde{S}(\hat{\vec{X}})$. Expanding $\tilde{S}(\hat{\vec{X}})$ in a Taylor series up to second order in $\Delta\hat{\vec{X}}$, we have
$\tilde{S}$ $(\vec{X}+\Delta\hat{\vec{X}})$
$=\tilde{S}(\vec{X})+\frac{\partial\tilde{S}(\vec{X})}{\partial\vec{X}}\cdot\Delta\hat{\vec{X}}+\frac{1}{2}\Delta\hat{\vec{X}}\cdot\frac{\partial^{2}\tilde{S}(\vec{X})}{\partial\vec{X}^{2}}\cdot\Delta\hat{\vec{X}}$
$=\tilde{S}(\vec{X})+\vec{Y}\cdot\Delta\hat{\vec{X}}-\frac{1}{2}\Delta\hat{\vec{X}}\cdot C^{-1}\cdot\Delta\hat{\vec{X}},$ (71)
where we used that $\frac{\partial\tilde{S}(\vec{X})}{\partial\vec{X}}=\vec{Y}$ and that the inverse of the fluctuation matrix (or covariance matrix) is $C^{-1}=-\frac{\partial\vec{Y}}{\partial\vec{X}}$. Replacing in (13), we obtain
$P(\Delta\hat{\vec{X}})=e^{-\tilde{S}[\vec{Y}]+\tilde{S}(\vec{X})-\vec{Y}\cdot\vec{X}-\frac{1}{2}\Delta\hat{\vec{X}}\cdot C^{-1}\cdot\Delta\hat{\vec{X}}}.$ (72)
Using the Legendre transform $S[\vec{Y}]=S(\vec{X})-\vec{Y}\cdot\vec{X}$, and $\Delta\tilde{S}(\vec{X})=\tilde{S}(\vec{X})-S(\vec{X})$ and $\Delta\tilde{S}[\vec{Y}]=\tilde{S}[\vec{Y}]-S[\vec{Y}]$, we have
$P(\Delta\hat{\vec{X}})=e^{-\Delta\tilde{S}[\vec{Y}]+\Delta\tilde{S}(\vec{X})}\,e^{-\frac{1}{2}\Delta\hat{\vec{X}}\cdot C^{-1}\cdot\Delta\hat{\vec{X}}},$ (73)
therefore, the factor $e^{-\Delta\tilde{S}[\vec{Y}]+\Delta\tilde{S}(\vec{X})}$, a function of the equilibrium parameters $\vec{X}$ and $\vec{Y}$, is the normalization of the multivariate normal distribution given by $(2\pi)^{-r/2}(\det C)^{-1/2}$, where $r$ is the number of variables; see, for example, Mishin (2015) or Sec. 7.C in Reichl (1998); the multivariate Gaussian distribution is a consequence of the central limit theorem, see Ch. V in Khinchin (1949).
@@
$e^{-\Delta\tilde{S}[\vec{Y}]+\Delta\tilde{S}(\vec{X})}=(2\pi)^{-r/2}(\det C)^{-1/2},$ (74)
@@
$\Delta\tilde{S}(\vec{X})-\Delta\tilde{S}[\vec{Y}]=-\frac{1}{2}\ln(\det C)-\frac{r}{2}\ln(2\pi).$ (75)
Now, using Eq. (37), we have that $\Delta\tilde{S}[\vec{Y}]$ is a constant:
$\Delta\tilde{S}[\vec{Y}]=c+\frac{r}{2}\ln(2\pi).$ (76)
@@
with $A=Me^{-2\epsilon/T}/\sqrt{N^{2}-M^{2}}$. Using this equation, the TH Massieu function $S[M,\tfrac{1}{T}]$ is
$S[M,$ $\tfrac{1}{T}]=\underbrace{N_{+}\ln\frac{N}{N_{+}}+N_{-}\ln\frac{N}{N_{-}}}_{\text{ideal part}}+\frac{\epsilon}{T}(2M-N)$
@@
And the mean squared fluctuations $\langle(\Delta\hat{M})^{2}\rangle=-(\frac{\partial^{2}S[M,1/T]}{\partial M^{2}})^{-1}$ are given by